OpenStudy (anonymous):

sin theta = sqrt 2/2 which could not be the value of theta? 225 degrees 45 degrees 135 degrees 405 degrees

OpenStudy (anonymous):



OpenStudy (anonymous):

Tell what cos and sin of 45 equal to ?

OpenStudy (anonymous):

1/sqrt2?

OpenStudy (anonymous):

yes, but you can't have a radical in the denominator.

OpenStudy (anonymous):

So, you would multiply top and bottom times √2, and you get, √2 / 2 .

OpenStudy (anonymous):

Ok, now, lets explore the other options. sin 225 = sin (180+45)=sin 180 cos 45 + cos 180 sin 45 = 0 (√2 / 2) + 1 (√2 / 2) = √2 / 2 So, option A works too. sin 135 = sin (180-45) = sin 180 cos 45 - cos 180 sin 45 = 0 (√2 / 2) - 1 (√2 / 2) = -√2 / 2 So, C doesn't work. sin 405 = sin (360 +45) = sin 360 cos 45 + cos 360 sin 45 = 0 (√2 / 2) + 1 (√2 / 2) =√2 / So option D works.. So options A B D work, option C doesn't. \(\LARGE\color{black}{ \bf :) }\)

OpenStudy (anonymous):

Thanks a ton! :D

OpenStudy (anonymous):

You welcome !